le f(x)= x^2-4x and g(x)=4/x-4, find (g0f)(-4)
Answer by stanbon(75887)

f(x)= x^2-4x and g(x)=4/(x-4),
find (gof)(-4)
= g[f(-4)
= g[16+16]
= g[32]
= 4/[32-4]
= 4/28
= 1/7
============
cheers,
Stan H.